# Optimization Insights into Deep Diagonal Linear Networks

### v1

@@ 2 Deep Diagonal Linear Networks
Figure 2: Representation of a Deep Diagonal Linear Network with $L$ layers in $\mathbb{R}^{d}$, $d=6$.
@@ 2.3 Convergence guarantees
Since we assume that assumption $\mathcal{A}$ holds, the second claim of Proposition 1 ensures that for any $i\in[d]$, if $j\notin\arg\min_{k\in[d]}|u^{k}_{i}(0)|$, then $|u^{j}_{i}(t)|>0$ for all $t\geqslant 0$. More precisely, using the first claim of Proposition 1, we obtain that $u^{j}_{i}(t)^{2}\geqslant u^{j}_{i}(0)^{2}-u^{l}_{i}(0)^{2}$ where $l=\arg\min_{k\in[d]}|u^{k}_{i}(0)|$. Therefore, for any $i\in[d]$,
@@
where $\sigma=\min_{i\in[d]}\prod_{k\neq k_{i}}(u^{k}_{i}(0)^{2}-u^{k_{i}}_{i}(0)^{2})$ and $k_{i}=\arg\min_{k\in[d]}|u^{k}_{i}(0)|$ for any $i\in[d]$.

### v2

@@ 2 Deep Diagonal Linear Networks
Figure 2: Representation of a Deep Diagonal Linear Network with $L$ layers in $\mathbb{R}^{d}$, $d=5$.
@@ 2.3 Convergence guarantees
Since we assume that assumption $\mathcal{A}$ holds, the second claim of Proposition 1 ensures that for any $i\in[d]$, if $j\notin\arg\min_{k\in[L]}|u^{k}_{i}(0)|$, then $|u^{j}_{i}(t)|>0$ for all $t\geqslant 0$. More precisely, using the first claim of Proposition 1, we obtain that $u^{j}_{i}(t)^{2}\geqslant u^{j}_{i}(0)^{2}-u^{l}_{i}(0)^{2}$ where $l=\arg\min_{k\in[L]}|u^{k}_{i}(0)|$. Therefore, for any $i\in[d]$,
@@
where $\sigma=\min_{i\in[d]}\prod_{k\neq k_{i}}(u^{k}_{i}(0)^{2}-u^{k_{i}}_{i}(0)^{2})$ and $k_{i}=\arg\min_{k\in[L]}|u^{k}_{i}(0)|$ for any $i\in[d]$.
@@ 3 Conclusion and Perspectives
Acknowledgements
H. L., L. R., C. M. and S. V. acknowledge the financial support of the European Research Council (grant SLING 819789), the European Commission (ELIAS 101120237), the US Air Force Office of Scientific Research (FA8655-22-1-7034), the Ministry of Education, University and Research (grant ML4IP R205T7J2KP; the European Commission (grant TraDE-OPT 861137); grant BAC FAIR PE00000013 funded by the EU - NGEU) and the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF-1231216. The research by C. M. and S. V. has been supported by the MUR Excellence Department Project awarded to Dipartimento di Matematica, Universita di Genova, CUP D33C23001110001. C. M. and S. V. are members of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains.